5.Balance the equation below and use it to answer the questions that follow:

C2O42-(aq) + MnO4-(aq) + H+(aq) = CO2(g) + Mn2+(aq) + H2O(I)

(I) Which ionic species is an oxidizing agent and which is a reducing agent

(II) Twenty five 25 cm3 of a solution of sodium ethanediote of concentration 0.1 mol dm-3 were placed in a conical flask. A burrete was filled with patassium manganate (VII) solution of concentration 0.038 mol dm-3 . Determine the volume of the manganate (VII) solution that would be needed to give the endpoint in the titration.

(III) Assuming that the KMnO4 was prepared by dissolving 6.94 g of salt in 1 dm3 of solution. Calculate the % purity of KMnO4

1
Expert's answer
2022-02-28T03:51:04-0500

C2O42-(aq) + MnO4-(aq) + 8H+(aq) = 2CO2(g) + Mn2+(aq) + 4H2O(I)

(I) 2MnVII + 10 e- → 2MnII (reduction).

10CIII - 10 e- → 10CIV (oxidation).

MnO4- is an oxidizing agent,

C2O42- is a reducing agent.

(II) C1V1 = C2V2

25 cm3 x 0.1 mol/dm3 = 0.030 mol/dm3 x V

V = 25 cm3 x 0.1 mol/dm3 / 0.030 mol/dm3 = 83.3 cm3

(III) n (KMnO4) = 6.94 g / 158 g/mol = 0.044 mol

% purity = 0.038 / 0.044 x 100% = 86.4%